Air flows steadily at the rate of 0.4 kg/s through an air compressor, entering at
6 m/s with a pressure of 100 kPa and a specific volume of 0.85 m3
/kg. The air
exits at 4.5 m/s with a pressure of 690 kPa and a specific volume of
0.16 m3
/kg. The specific internal energy of the air leaving is 88 kJ/kg greater
than that of the air entering. Cooling water in the jacket surrounding the
cylinder absorbs heat from the air at the rate of 59 kJ/s. Calculate:
a. The power required to drive the compressor
b. The cross-sectional area of the inlet pipe
c. The cross-sectional area of the outlet pipe
1
Expert's answer
2020-01-31T09:49:38-0500

"\dot m =0.4 \; kg/s"

"C_1 =6\; m/s, \quad p_1=100\times10^3 \;Pa, \quad v_1=0.85 \;m^3/kg"

"C_2 =4.5\; m/s, \quad p_2=690\times10^3 \;Pa, \quad v_2=0.16 \;m^3/kg"

"\Delta U=88\times 10^3 \; J/kg"

"\dot Q=-59\times 10^3 \; J/s"


a) "\dot W -?
\;\;"

b) "A_1 -?"

c) "A_2 -?"


Solution:

"C" - velocity, "p" - pressure, "v" - specific volume, "Z" - height, "U" - internal energy, "A" - cross-sectional area.

a) The steady flow energy equation:

"\dot Q-\dot W= \dot m\Delta H+\frac{1}{2} \dot m(C^2_2-C_1^2)+\dot mg(Z_2-Z_1)" (*)


"\Delta H=(U_2+ p_2v_2)-(U_1+p_1v_1)=\Delta U +p_2v_2-p_1v_1"

"\Delta H= 88\times 10^3+690\times10^3\times 0.16-100\times10^3\times0.85= 113.4\times 10^3 \;J/kg"


"\frac{1}2 \dot m(C_2^2-C_1^2)= \frac{1}2 \times0.4 \times (4.5^2-6^2)=-3.15\; J/kg"


"Z_1=Z_2"


(*): "-59\times 10^3 -\dot W=0.4\times 113.4\times 10^3-3.15"

"\dot W=-104.4 \;kW", negative sign indicates that the work is done on the system.


b,c) The continuity of mass equation:

"\dot m=\frac{C_1A_1}{V_1}=\frac{C_2A_2}{V_2}"


"A_1=\frac{\dot m V_1}{C_1}=0.0566 \; m^2"

"A_2=\frac{\dot m V_2}{C_2}=0.0142 \;m^2"


Answers:

a) "\dot W=-104.4 \; kW"

b) "A_1=0.057 \;m^2"

c) "A_2=0.014 \; m^2"